Q2. Consider the system described by the following differential equation x(t) Ax(t) whereand x(0) -
Answer #1

Given the differential equation is

i(t ) A2(t) ........ Eq.1

8 -6

r(0) =

Applying Laplace transform on Eq.1

sX(s)-X(0)=AX(s)

Rightarrow sIX(s)-AX(s)=X(0)

where I is a 2x2 identity matrix  01

Rightarrow (sI-A)X(s)=X(0)

X(s) = (sl-A)-1X(0) ........ Eq.2

-б

s+6 1

(sI-A)^{-1}=rac{1}{s^2+6s+8}egin{bmatrix} s+6 &1 -8& s end{bmatrix}

(sI-A)^{-1}=rac{1}{(s+2)(s+4)}egin{bmatrix} s+6 &1 -8& s end{bmatrix}...... Eq.3

From Eq.2 and Eq.3

X(s)=rac{1}{(s+2)(s+4)}egin{bmatrix} s+6 &1 -8& s end{bmatrix}egin{bmatrix} 1 1 end{bmatrix}

Rightarrow X(s)=rac{1}{(s+2)(s+4)}egin{bmatrix} s+7 s-8 end{bmatrix}

s+7 (s+2) (s+4)

Rightarrow X(s)=egin{bmatrix} rac{5}{2(s+2)}-rac{3}{2(s+4)} -rac{5}{(s+2)}+rac{6}{(s+4)} end{bmatrix}

Applying Inverse Laplace transform on the above equation,

-4t

Rightarrow x(t)=egin{bmatrix} 2.5e^{-2t}-1.5e^{-4t} --5e^{-2t}+6e^{-4t} end{bmatrix}

Solving i(t ) A2(t)

we get

x(t)=egin{bmatrix} 2.5e^{-2t}-1.5e^{-4t} --5e^{-2t}+6e^{-4t} end{bmatrix}